# Local Differential Privacy for Distributed Stochastic Aggregative Optimization with Guaranteed Optimality

### v2

@@ VII-A Auxiliary lemmas
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
×$\textstyle\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{\chi}^{k}\rangle]\leq\frac{\sqrt{d_{7}}d_{23}}{\lambda_{z}^{0}(t+1)^{\alpha_{2}}}$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
×$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{g}^{k}\rangle]\leq\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
×$\textstyle\times\sum_{k=2}^{t}(1-|\rho_{2}|)^{t-k}\frac{d_{26}}{(k+1)^{\beta_{1}+2v_{z}}}\leq\frac{d_{27}}{(t+1)^{\alpha_{4}}},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$ (120)
×$\textstyle\quad\!\times\!\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\!\otimes\!I_{r})^{p-k}\Theta_{g}^{k}\rangle]\!\leq\!\frac{\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{d_{28}}{\kappa(t+1)^{\beta_{1}}},$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
×$\textstyle\quad\times(\frac{1}{\lambda_{y}^{p}}(\boldsymbol{\hat{y}}^{p+1}-\boldsymbol{\hat{y}}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\boldsymbol{\hat{y}}^{p}-\boldsymbol{\hat{y}}^{p-1}))\rangle]$
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{25}}{(t+1)^{\alpha_{2}}},$ (121)
@@
$\textstyle\mathbb{E}\big{[}\big{\langle}\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}$ (122)
×$\textstyle\quad\times\lambda_{z}^{p}\left(\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p})-\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p-1})\right)\big{\rangle}\big{]}$
$\textstyle\leq\frac{3\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}+d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}+d_{25}}{(t+1)^{\alpha_{2}}}.$
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle]$ (123)
$\textstyle\leq\kappa\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{14}+d_{16}+d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}+d_{25}}{(t+1)^{\alpha_{2}}}.$
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\langle\bar{x}_{(i)}^{t}-x_{i}^{*},\nabla g_{i}^{t}(x_{i}^{t})\left(\frac{1}{m}\sum_{i=1}^{m}\tilde{z}_{i}^{t}-\tilde{z}_{i}^{t}\right)\rangle]$
$\textstyle\leq\frac{5\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{29}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{30}}{(t+1)^{\beta_{3}}},$ (124)
@@
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
×$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m\sigma_{g,1}^{2}L_{f}^{2}+mL_{g}^{2}\sigma_{f}^{2}}{\kappa(t+1)}$
@@
$\textstyle\leq\frac{(7\!+\!2(\sigma_{g,1}^{2}\!+\!L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{b_{3}}{\kappa(t+1)^{\beta_{1}}}\!+\!\frac{b_{4}}{(t+1)^{\beta_{3}}},$ (125)
@@
By substituting (61) and (83) into (60), we obtain
$\textstyle\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$ $\textstyle\leq\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@ VII-B Proof of Theorem 1
$\|\bar{x}_{(i)}^{t+1}-x_{i}^{*}\|^{2}\leq\|\bar{x}_{(i)}^{t}+\bar{\vartheta}_{w(i)}^{t}-\frac{\lambda_{x}^{t}}{m}u_{i}^{t}-x_{i}^{*}\|.$ (127)
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}-\frac{\lambda_{x}^{t}}{m}\boldsymbol{u}^{t}\|^{2}]+\mathbb{E}[\|\boldsymbol{\bar{\vartheta}}_{w}^{t}\|^{2}]$
$\textstyle\leq\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{(\lambda_{x}^{t})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t}\|^{2}]+n(\sigma_{x}^{t})^{2}$
$\textstyle\quad-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle].$ (128)
@@
$\vspace{-0.2em}\textstyle\frac{(\lambda_{x}^{t})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t}\|^{2}]\leq\frac{c_{2}}{(t+1)^{\varrho_{1}}},\vspace{-0.2em}$ (129)
where $\varrho_{1}=\min\{2v_{x}+2\varsigma_{y}-2v_{y},2v_{x}+2\varsigma_{z}-2v_{z}\}$ and $c_{2}=\frac{C_{z}c_{u1}(\lambda_{x}^{0})^{2}}{m^{2}(\lambda_{z}^{0})^{2}}+\frac{c_{u1}r(\lambda_{x}^{0})^{2}\sigma_{z}^{2}}{m(\lambda_{z}^{0})^{2}}+\frac{c_{u2}(\lambda_{x}^{0})^{2}}{m}$.
@@
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{\lambda_{x}^{t}\mu}{m}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$ (130)
$\textstyle\quad+\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle].$
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$ (131)
$\textstyle\leq\left(1-\frac{\lambda_{x}^{t}\mu}{m}+\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\lambda_{x}^{t}\kappa}{2m}\right)\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
$\vspace{-0.2em}\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}\!-\!\boldsymbol{x}^{*}\|^{2}]\!\leq\!(1-\frac{\lambda_{x}^{t}\mu}{2m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{c_{3}}{(t+1)^{\varrho_{2}}},$ (132)
@@
Using the relations $\varsigma_{y}>v_{y}-\frac{1}{2}$$\varsigma_{z}>v_{z}-\frac{1}{2}$, and $\varsigma_{x}>\frac{v_{x}}{2}$ from the theorem statement, we have $\varrho_{2}>v_{x}$. Therefore, according to Lemma 4 in [39], there exists a constant $c_{4}\triangleq\frac{\lambda_{x}^{0}\mu}{2mc_{3}}\max\{\mathbb{E}[\|\boldsymbol{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}],\frac{2mc_{3}}{\lambda_{x}^{0}\mu-2m(\varrho_{2}-v_{x})}\}$ such that the following inequality always holds:
$\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\leq\frac{2mc_{3}c_{4}}{\lambda_{x}^{0}\mu(t+1)^{\varrho_{2}-v_{x}}}=\frac{2mc_{3}c_{4}}{\lambda_{x}^{0}\mu(t+1)^{\beta}},$ (133)
@@
By using (50) and (133), we arrive at
$\textstyle\mathbb{E}[\|\boldsymbol{x}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq 2\sum_{i=1}^{m}\mathbb{E}[\|\boldsymbol{x}_{(i)}^{t}\!-\!\boldsymbol{1}_{m}\otimes\bar{x}_{(i)}^{t}\|^{2}]$ (134)
$\textstyle\quad+2\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\leq\frac{2C_{x}+4mc_{3}c_{4}(\lambda_{x}^{0}\mu^{-1})}{(t+1)^{\beta}},$
which implies (4) with $C_{x}$ given in (50), $c_{3}$ given in (132), and $c_{4}$ given in the paragraph above (133).
@@
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (135)
$\textstyle\quad-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{t},\boldsymbol{\breve{u}}^{t}\rangle]\!+\!\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle].$
The second term on the right hand side of (135) satisfies
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}\!-\!\boldsymbol{x}^{t},\boldsymbol{\breve{u}}^{t}\rangle]\leq\frac{2(L_{f}+L_{g}L_{f})\sqrt{C_{x}}\lambda_{x}^{0}}{m(t+1)^{\varsigma_{x}+v_{x}}}\triangleq\frac{c_{5}}{(t+1)^{\varsigma_{x}+v_{x}}},$
where in the derivation we have used (50).
Substituting the preceding inequality and (126) into (135) and letting $\kappa=\frac{2a_{t}}{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\lambda_{x}^{t}}$ for any $a_{t}>0$, we obtain
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (136)
$\textstyle\quad+\frac{a_{t}}{m}\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))b_{5}(\lambda_{x}^{t})^{2}}{a_{t}m(t+1)^{\beta_{1}}}$
@@
$\textstyle\mathbb{E}[\|\boldsymbol{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (137)
$\textstyle\quad+(1+\frac{a_{t}}{m})\mathbb{E}[\|\boldsymbol{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\Phi_{t},$

### v3

@@ VII-A Auxiliary lemmas
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
$\textstyle\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{\chi}^{k}\rangle]\leq\frac{\sqrt{d_{7}}d_{23}}{\lambda_{z}^{0}(t+1)^{\alpha_{2}}}$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$
$\textstyle\quad\times\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\otimes I_{r})^{p-k}\Theta_{g}^{k}\rangle]\leq\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
$\textstyle\times\sum_{k=2}^{t}(1-|\rho_{2}|)^{t-k}\frac{d_{26}}{(k+1)^{\beta_{1}+2v_{z}}}\leq\frac{d_{27}}{(t+1)^{\alpha_{4}}},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}M_{\tau^{p}}$ (114)
$\textstyle\quad\!\times\!\lambda_{z}^{p}\sum_{k=2}^{p}(\tilde{W}\!\otimes\!I_{r})^{p-k}\Theta_{g}^{k}\rangle]\!\leq\!\frac{\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{d_{28}}{\kappa(t+1)^{\beta_{1}}},$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\lambda_{z}^{p}M_{\tau^{p}}$
$\textstyle\quad\times(\frac{1}{\lambda_{y}^{p}}(\boldsymbol{\hat{y}}^{p+1}-\boldsymbol{\hat{y}}^{p})-\frac{1}{\lambda_{y}^{p-1}}(\boldsymbol{\hat{y}}^{p}-\boldsymbol{\hat{y}}^{p-1}))\rangle]$
$\textstyle\leq\frac{\kappa}{2}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{25}}{(t+1)^{\alpha_{2}}},$ (115)
@@
$\textstyle\mathbb{E}\big[\big\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}$ (116)
$\textstyle\quad\times\lambda_{z}^{p}\left(\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p})-\nabla_{y}\boldsymbol{f}(\boldsymbol{x}^{p-1},\boldsymbol{\tilde{y}}^{p-1})\right)\big\rangle\big]$
$\textstyle\leq\frac{3\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{16}+d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}+d_{25}}{(t+1)^{\alpha_{2}}}.$
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},-\frac{1}{\lambda_{z}^{t}}\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\sum_{p=1}^{t}(\tilde{W}\otimes I_{r})^{t-p}K_{2}\Xi_{f}^{p}\rangle]$ (117)
$\textstyle\leq\kappa\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{14}+d_{16}+d_{21}+d_{28}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{18}+d_{25}}{(t+1)^{\alpha_{2}}}.$
@@
$\textstyle\sum_{i=1}^{m}\mathbb{E}[\langle\bar{\mathrm{x}}_{(i)}^{t}-x_{i}^{*},\nabla g_{i}^{t}(x_{i}^{t})\left(\frac{1}{m}\sum_{i=1}^{m}\tilde{z}_{i}^{t}-\tilde{z}_{i}^{t}\right)\rangle]$
$\textstyle\leq\frac{5\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{d_{29}}{\kappa(t+1)^{\beta_{1}}}+\frac{d_{30}}{(t+1)^{\beta_{3}}},$ (118)
@@
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\nabla\boldsymbol{g}(\boldsymbol{x}^{t})$
$\textstyle\quad\times(\boldsymbol{1}_{m}\otimes\nabla_{y}\bar{f}(\boldsymbol{x}^{t},\boldsymbol{1}_{m}\otimes g(\boldsymbol{x}^{t}))-\nabla\boldsymbol{g}^{t}(\boldsymbol{x}^{t})\boldsymbol{\tilde{z}}^{t})\rangle]$
$\textstyle\leq\frac{(7+2(\sigma_{g,1}^{2}+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{m\sigma_{g,1}^{2}L_{f}^{2}+mL_{g}^{2}\sigma_{f}^{2}}{\kappa(t+1)}$
@@
$\textstyle\leq\frac{(7\!+\!2(\sigma_{g,1}^{2}\!+\!L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{b_{3}}{\kappa(t+1)^{\beta_{1}}}\!+\!\frac{b_{4}}{(t+1)^{\beta_{3}}},$ (119)
@@
By substituting (55) and ($7$) into (54), we obtain
$\textstyle\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle]$ $\textstyle\leq\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\kappa}{4}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@ VII-B Proof of Theorem 1
$\|\bar{\mathrm{x}}_{(i)}^{t+1}-x_{i}^{*}\|^{2}\leq\|\bar{\mathrm{x}}_{(i)}^{t}+\bar{\vartheta}_{w(i)}^{t}-\frac{\lambda_{x}^{t}}{m}u_{i}^{t}-x_{i}^{*}\|.$ (121)
@@
$\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{(\lambda_{x}^{t})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t}\|^{2}]$
$\textstyle\quad+n(\sigma_{x}^{t})^{2}-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle].$ (122)
@@
$\frac{(\lambda_{x}^{t})^{2}}{m^{2}}\mathbb{E}[\|\boldsymbol{u}^{t}\|^{2}]\leq\frac{c_{2}}{(t+1)^{\varrho_{1}}},$ (123)
where $\varrho_{1}=\min\{2v_{x}+2\varsigma_{y}-2v_{y},2v_{x}+2\varsigma_{z}-2v_{z}\}$ and $c_{2}=\frac{\bar{C}_{z}c_{u1}(\lambda_{x}^{0})^{2}}{m^{2}(\lambda_{z}^{0})^{2}}+\frac{c_{u1}r(\lambda_{x}^{0})^{2}\sigma_{z}^{2}}{m(\lambda_{z}^{0})^{2}}+\frac{c_{u2}(\lambda_{x}^{0})^{2}}{m}$.
@@
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{\lambda_{x}^{t}\mu}{m}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$ (124)
$\textstyle\quad+\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle].$
@@
$\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]$ (125)
$\textstyle\leq\left(1-\frac{\lambda_{x}^{t}\mu}{m}+\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\lambda_{x}^{t}\kappa}{2m}\right)\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]$
@@
$\vskip-1.99997pt\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}\!-\!\boldsymbol{x}^{*}\|^{2}]\!\leq\!(1-\frac{\lambda_{x}^{t}\mu}{2m})\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\!+\!\frac{c_{3}}{(t+1)^{\varrho_{2}}},$ (126)
@@
Using the relations $\varsigma_{y}>v_{y}-\frac{1}{2}$$\varsigma_{z}>v_{z}-\frac{1}{2}$, and $\varsigma_{x}>\frac{v_{x}}{2}$ from the theorem statement, we have $\varrho_{2}>v_{x}$. Therefore, according to Lemma 4 in [53], there exists a constant $c_{4}\triangleq\frac{\lambda_{x}^{0}\mu}{2mc_{3}}\max\{\mathbb{E}[\|\mathbf{\bar{x}}^{0}-\boldsymbol{x}^{*}\|^{2}],\frac{2mc_{3}}{\lambda_{x}^{0}\mu-2m(\varrho_{2}-v_{x})}\}$ such that the following inequality always holds:
$\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]\leq\frac{2mc_{3}c_{4}}{\lambda_{x}^{0}\mu(t+1)^{\varrho_{2}-v_{x}}}=\frac{2mc_{3}c_{4}}{\lambda_{x}^{0}\mu(t+1)^{\beta}},$ (127)
@@
$\textstyle\mathbb{E}[\|\mathbf{x}_{i}^{t}-\boldsymbol{x}^{*}\|^{2}]\leq 2\mathbb{E}[\|\mathbf{x}_{i}^{t}\!-\!\mathbf{\bar{x}}^{t}\|^{2}]+2\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]$ (128)
$\textstyle\leq 2\sum_{i=1}^{m}\mathbb{E}[\|\mathbf{x}_{(i)}^{t}-\boldsymbol{1}_{m}\otimes\bar{\mathrm{x}}_{(i)}^{t}\|^{2}]+2\mathbb{E}[\|\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{*}\|^{2}]$
$\textstyle\leq\frac{2\bar{C}_{x}+4mc_{3}c_{4}(\lambda_{x}^{0}\mu^{-1})}{(t+1)^{\beta}},$
which implies (5) with $\bar{C}_{x}$ given in (44), $c_{3}$ given in (126), and $c_{4}$ given in the paragraph above (127).
@@
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (129)
$\textstyle\quad-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}\!-\!\boldsymbol{x}^{t},\boldsymbol{\breve{u}}^{t}\rangle]\!+\!\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{\breve{u}}^{t}-\boldsymbol{u}^{t}\rangle].$
Substituting (44) and (120) into (129) and letting $\kappa=\frac{2a_{t}}{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))\lambda_{x}^{t}}$ for any $a_{t}>0$, we obtain
$\textstyle-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[\langle\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*},\boldsymbol{u}^{t}\rangle]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (130)
$\textstyle\quad+\frac{a_{t}}{m}\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\frac{(9+2(\sigma_{g,1}^{2}\!+L_{g}^{2}))b_{5}(\lambda_{x}^{t})^{2}}{a_{t}m(t+1)^{\beta_{1}}}$
@@
$\textstyle\mathbb{E}[\|\mathbf{\bar{x}}^{t+1}-\boldsymbol{x}^{*}\|^{2}]\leq-\frac{2\lambda_{x}^{t}}{m}\mathbb{E}[F(\boldsymbol{x}^{t})-F(\boldsymbol{x}^{*})]$ (131)
$\textstyle\quad+(1+\frac{a_{t}}{m})\mathbb{E}[\|\mathbf{\bar{x}}^{t}-\boldsymbol{x}^{*}\|^{2}]+\Phi_{t},$